how can u graph x-2y+5 by finding the slope
Found 2 solutions by mananth, Alan3354:
Answer by mananth(16946)

x-2y+5=0
2y=x+5
y=(1/2)x +5/2
slope = 1/2


x=3, y=4 (3,4)
x=-3 y= 1 (-3,1)
x=9,y=7 (9,7)
your graph will look as below
.

Answer by Alan3354(69443)

how can u graph x-2y+5 by finding the slope
----------
I can't, because it's not an equation.
I know that, because there's no equal sign.